A wheel made 374 revolutions at 40 7/8 revolutions per minute. How many minutes did it rotate?
Answer by Ave(106)

you can use distance=rate*time for this
374=40 7/8t first convert the mixed number it into an improper fraction
320/8+7/8=327/8
374=327/8t multiply by 8 to get rid of fraction
2992=327t
9.1=t
9.1minutes